Let a line passing through (2, 1) intersects the x-axis at A and y-axis at B. Find the locus of the midpoint of AB.

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
let mid point is (h,k)
now 
 
x/a + y/b = 1
 
2/a + 1/b = 1
 
also a/2 = h and b/2 = k
 
hence
 
2/2h + 1/2k = 1